# Rigidity of anti-de Sitter (2+1)-spacetimes with convex boundary near the Fuchsian locus

### v1

@@ 4.4. Proof of the main result
Both spaces $\mathcal{E}$ and $\mathcal{S}$ are nuclear, hence paracompact [38, Theorem 16.10]. Since they also are Hausdorff and locally metrizable, they are metrizable by the Smirnov metrization theorem [45, Theorem 42.1]. Choose a metric $d_{S}$ on $S$ and consider its restriction to $\mathcal{S}_{F}:=\mathcal{I}(\mathcal{E}_{F})$. Because by Labourie–Schlenker [41], $\mathcal{I}|_{\mathcal{E}_{F}}$ is a homeomorphism onto the image, we can pullback $d_{S}$ from $\mathcal{S}_{F}$ to $\mathcal{E}_{F}$. The subset $\mathcal{E}_{F}$ is closed in $\mathcal{E}$ and it is a result of Hausdorff [35] that a metric on $\mathcal{E}_{F}$ extends to a metric $d_{E}$ defining the topology of $\mathcal{E}$.
Now from Lemma 4.5 for every $e\in\mathcal{E}_{F}$ there is $r(e)>0$ such that $\mathcal{I}$ is a homeomorphism onto the image when restricted to the open ball $B(e,r(e))$. We can pick a neighborhood $U_{e}$ of $e$ such that $U_{e}\subset B(e,r(e)/3)$ and $\mathcal{I}(U_{e})\subset B(\mathcal{I}(e),r(e)/3)$. We claim that $\mathcal{I}$ is injective on $U=\bigcup_{e\in\mathcal{E}_{F}}U_{e}$, which is then a desired neighborhood of $\mathcal{E}_{F}$. Indeed, suppose that $h=\mathcal{I}(e_{1}^{\prime})=\mathcal{I}(e_{2}^{\prime})$ for $e_{1}^{\prime},e_{2}^{\prime}\in U$, $e_{1}^{\prime}\in U_{e_{1}}$, $e_{2}^{\prime}\in U_{e_{2}}$ with $e_{1},e_{2}\in\mathcal{E}_{F}$, and that $r_{1}:=r(e_{1})\leq r_{2}:=r(e_{2})$. We claim that then $e_{1}^{\prime}\in U_{e_{2}}$. Indeed,
@@ 5.1. Anti-de Sitter spacetimes with marked points
Every GHC AdS (2+1)-spacetime $M$ admits a canonical extension to a GHMC AdS (2+1)-spacetime, which we denote for now by $N$. The spacetime $N$ has a convex core, which can be defined, e.g., as the inclusion-minimal closed totally convex subset of $N$. See details, e.g., in [15]. In particular, the convex core is contained in $M$ (considered here as a subset of $N$). We say that $M$ has strictly polyhedral boundary if it is locally modeled on convex polyhedral subsets of $\mathbb{A}d\mathbb{S}^{3}$ and the boundary does not intersect the convex core. The last condition ensures a reasonable behavior of the combinatorics of the boundary as we vary the metric on $M$. See [50, Section 3] on the description of this in the hyperbolic case, as the anti-de Sitter case is the same in this regard. By a vertex of $M$ we mean a point on $\partial M$ at which $M$ is locally modeled on a convex cone different from a half-space or from the wedge of two half-spaces.
@@
The space $\mathscr{M}(V)$ is endowed with the topology of uniform convergence of developing maps on compact subsets of $\widetilde{M}$. As well as in the smooth case, we actually care only about the structure on the boundary. However, in this case the objects should be thought in a slightly different fashion.
Pick $\rho\in\widehat{\mathcal{R}}$, recall that $D(\rho)\subset\mathbb{A}d\mathbb{S}^{3}$ is the maximal open $\rho$-invariant convex set, on which the action of $\rho$ is free and properly discontinuous. It also contains the convex core $C(\rho)$, which is the inclusion-minimal closed $\rho$-invariant convex set. Alternatively, it can be defined as the convex hull of the limit set $\Lambda(\rho)$ at the boundary at infinity of $\mathbb{A}d\mathbb{S}^{3}$. (Recall that the $\rho$-quotient of $D(\rho)$ is a GHMC AdS spacetime. Then the convex core of this spacetime, mentioned above, is exactly the quotient of $C(\rho)$.) There are two connected components of $D(\rho)\backslash C(\rho)$, corresponding to the future-convex and past-convex components of $\partial D(\rho)$. We denote these components of $D(\rho)\backslash C(\rho)$ by $D_{+}(\rho)$ and $D_{-}(\rho)$.
@@
For now, let $V\subset S$ be a finite subset and $\widetilde{V}\subset\widetilde{S}$ be its full preimage, equipped with the natural $\pi_{1}S$ action. Let $\rho\in\widehat{\mathcal{R}}$ and let $f:\widetilde{V}\rightarrow D_{+}(\rho)$ be a $\rho$-equivariant map. Denote by $\overline{\rm conv}(f)$ the closed convex hull of $f(\widetilde{V})$ in $\mathbb{A}d\mathbb{S}^{3}$. Only one of its boundary components has nonempty intersection with $D_{+}(\rho)$. This component is spacelike and future-convex. We denote it by $\Sigma(f)$. We say that $f$ is future-convex if no point of $f(\widetilde{V})$ is a convex combination of others and strictly future-convex if, in addition, $\Sigma(f)$ is disjoint from $C(\rho)$. Note that then $\Sigma(f)\subset D_{+}(\rho)$, $f(\widetilde{V})\subset\Sigma(f)$ and $\Sigma(f)$ has a a decomposition into vertices, edges and faces, where the set of vertices is exactly $f(\widetilde{V})$, every edge is a geodesic segment between two vertices, and every face is isometric to a convex hyperbolic polygon. For more details on this decomposition, we refer to [50, Section 3], describing the situation in the hyperbolic case, which translates directly to the anti-de Sitter situation.
@@
Denote the space of quasi-Fuchsian equivariant strictly future-convex maps by ^$\widehat{\mathscr{E}}_{+}(V)$. By picking a lift in $\widetilde{V}$ for each $v\in V$, the space ^$\widehat{\mathscr{E}}_{+}(V)$ can be considered as a subset of ^×$\widehat{\mathcal{R}}\times(\mathbb{A}d\mathbb{S}^{3})^{V}$, from which it inherits a topology. This is an open subset, hence $\widehat{\mathscr{E}}_{+}$ becomes a smooth manifold. Denote by $\mathscr{E}_{+}(V)$ its quotient by $G$. It is a smooth manifold of dimension $12k-12+3n$, where $k$ is the genus of $S$ and $n:=|V|$. Similarly, we define quasi-Fuchsian equivariant strictly past-convex maps and denote the respective spaces by ^$\widehat{\mathscr{E}}_{-}(V)$ and $\mathscr{E}_{-}(V)$.
@@
A tangent vector to ^$\widehat{\mathscr{E}}_{+}(V)$ at $(\rho,f)$ is a vector field $v$ on $f(\widetilde{V})$ satisfying the automorphicity condition (2.3). Again, we say that $v$ is trivial if it is the restriction of a global Killing field. Pick an arbitrary triangulation of $\Sigma(f)$ subdividing its face decomposition. We say that $v$ is isometric if it induces zero variation on each edge-length of the triangulation. Note that this is independent on the chosen triangulation.
Let $\mathscr{S}(V)$ be the space of isotopy classes of hyperbolic cone-metrics on $S$ with cone-angles $>2\pi$ with vertices at $V$. Here we mean that every point of $V$ has cone-angle $>2\pi$ in such metrics. Every such metric admits a geodesic triangulation and $\mathscr{S}(V)$ is endowed with a topology of a smooth manifold of dimension $12k-12+3n$ by considering as charts the edge-length maps for varying triangulations. See, e.g., [49, Section 2.2] for details.
Return to the space $\mathscr{M}(V)$. Denote by $V_{+}$ and $V_{-}$ the intersections of $V$ with both boundary components of $M$, considered as subsets of $S$ via an identification $M\cong S\times[-1,1]$, and redefine $V:=\{V_{+},V_{-}\}$. Denote by ^$\widehat{\mathscr{E}}(V)$ the subset of ^$\widehat{\mathscr{E}}_{+}(V_{+})\times\widehat{\mathscr{E}}_{-}(V_{-})$ consisting of the pairs of maps that are equivariant with respect to the same representation, and by $\mathscr{E}(V)$ denote its quotient by $G$. We have a natural continuous injective map $\mathscr{M}(V)\hookrightarrow\mathscr{E}(V)$. Imitating the proof from the smooth case, we obtain that this map is a homeomorphism. We denote by $\mathscr{E}_{F}(V)$ the subspace of Fuchsian pairs of maps, i.e., the ones that have a lift in ^$\widehat{\mathscr{E}}_{F}$ with representation into $G_{F}$.
Pick $e\in\mathscr{E}(V)$ represented by $(\rho,f_{+},f_{-})\in\widehat{\mathscr{E}}(V)$. We have two associated $\rho$-invariant spacelike surfaces $\Sigma(f_{+})$ and $\Sigma(f_{-})$. They are endowed with hyperbolic cone-metrics with cone-angles $>2\pi$. Consider their $\rho$-quotients. We have a canonical isomorphism of the fundamental group of each of these quotients with $\pi_{1}S$. This allows to define homeomorphisms from them to $S$, determined up to isotopy. Furthermore, the markings of vertices given by $f_{+}$ and $f_{-}$ allow to associate the induced metrics with elements of $\mathscr{S}(V_{+})$ and $\mathscr{S}(V_{-})$ respectively. This defines the induced metric map
@@
There exists a neighborhood $U$ of $\mathscr{E}_{F}(V)$ in $\mathscr{E}(V)$ and an open subset $U^{\prime}\subset\mathscr{S}(V_{+})\times\mathscr{S}(V_{-})$ such that $\mathscr{I}_{V}|_{U}$ is a homeomorphism onto $U^{\prime}$
@@
The proof is identical to the proof of respective results in other settings, see, e.g., [25, Lemma 2.13] for a Minkowski version, or [50, Lemma 2.17] for a de Sitter version.
Lemma 5.4.
@@
Because $\mathscr{E}(V)$ and $\mathscr{S}(V_{+})\times\mathscr{S}(V_{-})$ are finite-dimensional manifolds of the same dimension, Lemma 5.4 actually means that for each $e\in\mathscr{E}_{F}(V)$, $d_{e}\mathscr{I}_{V}$ is an isomorphism. Together with Lemma 5.3, the inverse function theorem implies that $\mathscr{I}_{V}$ is a local homeomorphism around $\mathscr{E}_{F}(V)$. Using Theorem 5.5 instead of Theorem 2.3, the proof is finished in the same way as the proof of Theorem 2.1. ∎
@@ 5.2. Minkowski spacetimes with marked points and infinitesimal rigidity
In this section we prove Lemma 5.4. As in the smooth case, the proof is done by using the Pogorelov map. We will rely on [25] by Fillastre–Prosanov instead of [58]. We now recall the basics of polyhedral surfaces in Minkowski 3-space $\mathbb{R}^{2,1}$. We adapt some notions from the previous subsection to this setting.
Pick $\rho\in\widehat{\mathcal{R}}^{0}$. Associated with it are two maximal open $\rho$-invariant convex sets, on which the action of $\rho$ is free and properly discontinuous, one is future-complete and one is past-complete. We denote them by $D_{+}(\rho)$ and $D_{-}(\rho)$. See [44] for details.
Again, let $V\subset S$ be a finite subset and $\widetilde{V}\subset\widetilde{S}$ be its full preimage, equipped with the natural $\pi_{1}S$ action. Let $\rho\in\widehat{\mathcal{R}}^{0}$ and let $f:\widetilde{V}\rightarrow D_{+}(\rho)$ be a $\rho$-equivariant map. Denote by $\overline{\rm conv}(f)$ the closed convex hull of $f(\widetilde{V})$. In this setting, it is a future-complete set with one boundary component. Furthermore, this component is future-convex and spacelike. We denote it by $\Sigma(f)$. We say that $f$ is future-convex if no point of $f(\widetilde{V})$ is a convex combination of others. (An interesting difference with the anti-de Sitter situation is that here is no notion of strict convexity for our maps due to the lack of convex cores in this setting.) Note that then $f(\widetilde{V})\subset\Sigma(f)$. The surface $\Sigma(f)$ also has a a decomposition into vertices, edges and faces, where the set of vertices is exactly $f(\widetilde{V})$, every edge is a geodesic segment between two vertices, and every face is isometric to a convex Euclidean polygon.
Denote the space of quasi-Fuchsian equivariant future-convex maps by ^$\widehat{\mathscr{E}}_{+}^{0}(V)$. By picking a lift in $\widetilde{V}$ for each $v\in V$, the space ^$\widehat{\mathscr{E}}_{+}(V)$ can be considered as a subset of ^$\widehat{\mathcal{R}}^{0}\times(\mathbb{R}^{2,1})^{V}$, from which it inherits a topology. This is an open subset, hence ^$\widehat{\mathscr{E}}_{+}^{0}(V)$ becomes a smooth manifold. Denote by $\mathscr{E}_{+}^{0}(V)$ its quotient by $G$. It is a smooth manifold of dimension $12k-12+3n$, where $k$ is the genus of $S$ and $n:=|V|$. Similarly, we define quasi-Fuchsian equivariant past-convex maps and denote the respective spaces by ^$\widehat{\mathscr{E}}_{-}^{0}(V)$ and $\mathscr{E}_{-}^{0}(V)$. The notions of automorphic, trivial and isometric vector fields extend to this setting accordingly.
Let $\mathscr{S}^{0}(V)$ be the space of Euclidean cone-metrics with cone-angles $>2\pi$ and vertices at $V$, considered up to isotopy. For $e\in\mathscr{E}_{+}^{0}(V)$ lifting to $(\rho,f)$ the induced metric on $\Sigma(f)$ can be associated with an element in $\mathscr{S}^{0}(V)$. This defines the induced metric map
@@
It was shown in [25, Lemma 2.13] that this map is $C^{1}$. Moreover, the authors proved
@@
This is a direct corollary of [25, Theorem I]. See also the discussion in [25, Section 4.1]. For $s\in\mathscr{S}^{0}(V)$ denote $(\mathscr{I}_{V,+}^{0})^{-1}(s)$ by $\mathscr{E}_{+}^{0}(s)$. Due to Theorem 5.6, this is a $C^{1}$-submanifold of $\mathscr{E}_{+}^{0}(V)$. We now define a map
@@
which sends a $\rho$-equivariant map to $\rho$ (recall that $\mathcal{R}^{0}\cong T\mathscr{T}$). It is also $C^{1}$, see [25, Section 4.1]. We define similarly the space $\mathscr{E}^{0}_{-}(s)$ and the map $\phi^{0}_{s,-}$. It is another direct corollary of [25, Theorem I] that
@@
For any Euclidean cone-metric $s$ with cone-angles $>2\pi$ the compositions of $\phi_{s,\pm}^{0}$ with the projection to $\mathscr{T}$ are $C^{1}$-diffeomorphisms.
In particular, the maps $\phi^{0}_{s,\pm}$ are $C^{1}$-immersions. From [25, Theorem II], we have
@@
Pick two finite subsets $V_{+},V_{-}\subset S$ and metrics $s_{+}\in\mathscr{S}^{0}(V_{+})$, $s_{-}\in\mathscr{S}^{0}(V_{-})$. Then the intersection of ϕs+,+0subscriptsuperscriptitalic-$\phi^{0}_{s_{+},+}$ and ϕs,0subscriptsuperscriptitalic-$\phi^{0}_{s_{-},-}$ is unique and transverse.
@@
Pick two finite subsets $V_{+},V_{-}\subset S$. Consider $e_{+}=(\rho,f_{+})\in\mathscr{E}_{+}^{0}(V_{+})$, $e_{-}=(\rho,f_{-})\in\mathscr{E}_{-}^{0}(V_{-})$ and $\dot{\rho}\in Z^{1}_{\rho}(\pi_{1}S,\mathfrak{g}^{0})$. Let $v$ be a $\dot{\rho}$-automorphic isometric vector field on $f_{+}(\widetilde{V}_{+})$ and $f_{-}(\widetilde{V}_{-})$. Then $v$ is trivial.
@@
We can now prove Lemma 5.4.
Proof of Lemma 5.4..
As in the proof of Lemma 3.1, we exploit the simultaneous model for anti-de Sitter and Minkowski 3-spaces as well as the Pogorelov map. Recall the model from Section 3.1, particularly the point $o$ and the set $C$. Pick $e\in\mathscr{E}_{F}(V)$ and lift it to a triple $(\rho,f_{+},f_{-})\in\widehat{\mathscr{E}}_{F}(V)$, where $\rho:\pi_{1}S\rightarrow G_{F}$ and $G_{F}$ fixes $o$. Then the images of $f_{+}$ and $f_{-}$ are contained in $C$. Pick $\dot{e}\in T_{e}\mathscr{E}(V)$ such that $d_{e}\mathscr{I}_{V}(\dot{e})=0$. Then it is represented by a $\dot{\rho}$-automorphic isometric vector field on $f_{+}(\widetilde{V}_{+})$ and $f_{-}(\widetilde{V}_{-})$ for $\dot{\rho}\in Z^{1}_{\rho}(\pi_{1}S,\mathfrak{g})$.
Recall the Pogorelov map $\Phi$ and the associated map $\Psi_{\rho}$. From Lemma 3.11, $\Phi(v)$ is a $\Psi_{\rho}(\dot{\rho})$-automorphic vector field. Also, if $v$ induces a zero variation on the distance between two points, then this means that its restriction to these two points coincides with the restriction of a Killing field. Hence, Lemma 3.8 implies that $\Phi(v)$ is isometric. By Lemma 5.9, it is trivial, thereby so is $v$. It follows that $\dot{e}=0$. ∎

### v2

@@ 4.4. Proof of the main result
Both spaces $\mathcal{E}$ and $\mathcal{S}\times\mathcal{S}$ are nuclear, hence paracompact [40, Theorem 16.10]. Since they also are Hausdorff and locally metrizable, they are metrizable by the Smirnov metrization theorem [47, Theorem 42.1]. Choose a metric $d_{S}$ on $\mathcal{S}\times\mathcal{S}$ and consider its restriction to $\mathcal{S}_{F}:=\mathcal{I}(\mathcal{E}_{F})$. By Labourie–Schlenker [43], $\mathcal{I}|_{\mathcal{E}_{F}}$ is a homeomorphism onto the image, we can pullback $d_{S}$ from $\mathcal{S}_{F}$ to $\mathcal{E}_{F}$. The subset $\mathcal{E}_{F}$ is closed in $\mathcal{E}$ and it is a result of Hausdorff [37] that a metric on $\mathcal{E}_{F}$ extends to a metric $d_{E}$ on $\mathcal{E}$ compatible with the topology of $\mathcal{E}$.
Now from Proposition 4.5, for every $e\in\mathcal{E}_{F}$ there is $r(e)>0$ such that $\mathcal{I}$ is a homeomorphism onto the image when restricted to the closed ball $B(e,r(e))$. We can pick a neighborhood $U_{e}$ of $e$ such that $U_{e}\subset B(e,r(e)/3)$ and $\mathcal{I}(U_{e})\subset B(\mathcal{I}(e),r(e)/3)$. We claim that $\mathcal{I}$ is injective on $U=\bigcup_{e\in\mathcal{E}_{F}}U_{e}$, which is then a desired neighborhood of $\mathcal{E}_{F}$. Indeed, suppose that $h=\mathcal{I}(e_{1}^{\prime})=\mathcal{I}(e_{2}^{\prime})$ for $e_{1}^{\prime},e_{2}^{\prime}\in U$, $e_{1}^{\prime}\in U_{e_{1}}$, $e_{2}^{\prime}\in U_{e_{2}}$ with $e_{1},e_{2}\in\mathcal{E}_{F}$, and that $r_{1}:=r(e_{1})\leq r_{2}:=r(e_{2})$. We claim that then $e_{1}^{\prime}\in B(e_{2},r(e_{2}))$. Indeed,
@@ 5.1. Anti-de Sitter spacetimes with marked points
Every GHC AdS (2+1)-spacetime $M$ admits a canonical extension to a GHMC AdS (2+1)-spacetime, which we denote for now by $N$. The spacetime $N$ has a convex core, which can be defined, e.g., as the inclusion-minimal closed totally convex subset of $N$. See details, e.g., in [17]. In particular, the convex core is contained in $M$ (considered here as a subset of $N$). We say that $M$ has strictly polyhedral boundary if it is locally modeled on convex polyhedral subsets of $\mathbb{A}d\mathbb{S}^{3}$ and the boundary does not intersect the convex core. The last condition ensures a reasonable behavior of the combinatorics of the boundary as we vary the metric on $M$. See [51, Section 3] on the description of this in the hyperbolic case, as the anti-de Sitter case is the same in this regard. By a vertex of $M$ we mean a point on $\partial M$ at which $M$ is locally modeled on a convex cone different from a half-space or from the wedge of two half-spaces.
@@
The space $\mathscr{M}(V)$, defined in the introduction, is endowed with the topology of uniform convergence of developing maps on compact subsets of $\widetilde{M}$. As well as in the smooth case, we actually care only about the structures on the boundary. However, in this case the objects should be thought in a slightly different fashion.
[…]
@@
For now, let $V\subset S$ be a finite subset and $\widetilde{V}\subset\widetilde{S}$ be its full preimage, equipped with the natural $\pi_{1}S$ action. Let $\rho\in\widehat{\mathcal{R}}$ and let $f:\widetilde{V}\rightarrow D_{+}(\rho)$ be a $\rho$-equivariant map. Denote by $\overline{\rm conv}(f)$ the closed convex hull of f(V~)f($\widetilde${V}) in $\mathbb{A}d\mathbb{S}^{3}$. Only one of its boundary components has nonempty intersection with $D_{+}(\rho)$. This component is spacelike and future-convex. We denote it by $\Sigma(f)$. We say that $f$ is future-convex if no point of f(V~)f($\widetilde${V}) is a convex combination of others and strictly future-convex if, in addition, $\Sigma(f)$ is disjoint from $C(\rho)$. Note that then $\Sigma(f)\subset D_{+}(\rho)$, $f(\widetilde{V})\subset\Sigma(f)$ and $\Sigma(f)$ has a a decomposition into vertices, edges and faces, where the set of vertices is exactly f(V~)f($\widetilde${V}), every edge is a geodesic segment between two vertices, and every face is isometric to a convex hyperbolic polygon. For more details on this decomposition, we refer to [54, Section 3.4.5].
@@
Denote the space of quasi-Fuchsian equivariant strictly future-convex maps by $\widehat{\mathscr{E}}_{+}(V)$. By picking a lift in $\widetilde{V}$ for each $v\in V$, the space $\widehat{\mathscr{E}}_{+}(V)$ can be considered as a subset of $\widehat{\mathcal{R}}\times(\mathbb{A}d\mathbb{S}^{3})^{V}$, from which it inherits a topology. This is an open subset, hence $\widehat{\mathscr{E}}_{+}$ becomes a smooth manifold. Denote by $\mathscr{E}_{+}(V)$ its quotient by $G$. It is a smooth manifold of dimension 6k6+3n6k-6+3n, where $k$ is the genus of $S$ and $n:=|V|$. Similarly, we define quasi-Fuchsian equivariant strictly past-convex maps and denote the respective spaces by $\widehat{\mathscr{E}}_{-}(V)$ and $\mathscr{E}_{-}(V)$.
@@
A tangent vector to $\widehat{\mathscr{E}}_{+}(V)$ at $(\rho,f)$ is a vector field $v$ on f(V~)f($\widetilde${V}) satisfying the automorphicity condition (2.3). Again, we say that $v$ is trivial if it is the restriction of a global Killing field. Pick an arbitrary triangulation of $\Sigma(f)$ subdividing its face decomposition. We say that $v$ is isometric if it induces zero variation on each edge-length of the triangulation. Note that this is independent on the chosen triangulation.
Let $\mathscr{S}(V)$ be the space of isotopy classes of hyperbolic cone-metrics on $S$ with cone-angles $2\pi$ with vertices at $V$. Here we mean that every point of $V$ has cone-angle $2\pi$ in such metrics. Every such metric admits a geodesic triangulation and $\mathscr{S}(V)$ is endowed with a topology of a smooth manifold of dimension 6k6+3n6k-6+3n by considering as charts the edge-length maps for varying triangulations. See, e.g., [53, Section 2.2] for details.
Return to the space $\mathscr{M}(V)$ defined in the introduction. Denote by $V_{+}$ and $V_{-}$ the intersections of $V$ with both boundary components of $M$, considered as subsets of $S$ via an identification $M\cong S\times[-1,1]$, and redefine $V:=\{V_{+},V_{-}\}$. Denote by $\widehat{\mathscr{E}}(V)$ the subset of $\widehat{\mathscr{E}}_{+}(V_{+})\times\widehat{\mathscr{E}}_{-}(V_{-})$ consisting of the pairs of maps that are equivariant with respect to the same representation, and by $\mathscr{E}(V)$ denote its quotient by $G$. We have a natural continuous injective map $\mathscr{M}(V)\hookrightarrow\mathscr{E}(V)$. Imitating the proof from the smooth case, we obtain that this map is a homeomorphism. We denote by $\mathscr{E}_{F}(V)$ the subspace of Fuchsian pairs of maps, i.e., the ones that have a lift in $\widehat{\mathscr{E}}_{F}$ with representation into $G_{F}$.
Pick $e\in\mathscr{E}(V)$ represented by $(\rho,f_{+},f_{-})\in\widehat{\mathscr{E}}(V)$. We have two associated $\rho$-invariant spacelike surfaces $\Sigma(f_{+})$ and $\Sigma(f_{-})$. They are endowed with hyperbolic cone-metrics with cone-angles $2\pi$. Consider their $\rho$-quotients. We have a canonical isomorphism of the fundamental group of each of these quotients with $\pi_{1}S$. This allows to define homeomorphisms from them to $S$, determined up to isotopy. Furthermore, the markings of vertices given by $f_{+}$ and $f_{-}$ allow to associate the induced metrics with elements of $\mathscr{S}(V_{+})$ and $\mathscr{S}(V_{-})$ respectively. This defines the induced metric map
@@
There exists a neighborhood $U$ of $\mathscr{E}_{F}(V)$ in $\mathscr{E}(V)$ and an open subset $U^{\prime}\subset\mathscr{S}(V_{+})\times\mathscr{S}(V_{-})$, which contains the diagonal, such that $\mathscr{I}_{V}|_{U}$ is a homeomorphism onto $U^{\prime}$
@@
The proof is identical to the proof of respective results in other settings, see, e.g., [31, Lemma 2.13] for a Minkowski version or [51, Lemma 2.17] for a de Sitter version.
Proposition 5.4.
@@
Because $\mathscr{E}(V)$ and $\mathscr{S}(V_{+})\times\mathscr{S}(V_{-})$ are finite-dimensional manifolds of the same dimension, Proposition 5.4 actually means that for each $e\in\mathscr{E}_{F}(V)$, $d_{e}\mathscr{I}_{V}$ is an isomorphism. Together with Lemma 5.3, the inverse function theorem implies that $\mathscr{I}_{V}$ is a local homeomorphism around $\mathscr{E}_{F}(V)$. Using Theorem 5.5 instead of Theorem 2.4, the proof is finished in the same way as the proof of Theorem 2.1. ∎
@@ 5.2. Minkowski spacetimes with marked points and infinitesimal rigidity
In this section we prove Proposition 5.4. As in the smooth case, the proof is done by using the Pogorelov map. We will rely on [31] by Fillastre–Prosanov instead of [62]. We now recall the basics of polyhedral surfaces in Minkowski 3-space $\mathbb{R}^{2,1}$. We adapt some notions from the previous subsection to this setting.
Pick $\rho\in\widehat{\mathcal{R}}^{0}$. Associated with it are two maximal open $\rho$-invariant convex sets, on which the action of $\rho$ is free and properly discontinuous, one is future-complete and one is past-complete. We denote them by $D_{+}^{0}(\rho)$ and $D_{-}^{0}(\rho)$. See [46] of Mess for details.
Again, let $V\subset S$ be a finite subset and $\widetilde{V}\subset\widetilde{S}$ be its full preimage, equipped with the natural $\pi_{1}S$ action. Let $\rho\in\widehat{\mathcal{R}}^{0}$ and let $f:\widetilde{V}\rightarrow D_{+}^{0}(\rho)$ be a $\rho$-equivariant map. Denote by $\overline{\rm conv}(f)$ the closed convex hull of f(V~)f($\widetilde${V}). In this setting, it is a future-complete set with one boundary component. Furthermore, this component is future-convex and spacelike. We denote it by $\Sigma(f)$. We say that $f$ is future-convex if no point of f(V~)f($\widetilde${V}) is a convex combination of others. (An interesting difference with the anti-de Sitter situation is that here is no notion of strict convexity for our maps due to the lack of convex cores in this setting.) Note that then $f(\widetilde{V})\subset\Sigma(f)$. The surface $\Sigma(f)$ also has a a decomposition into vertices, edges and faces, where the set of vertices is exactly f(V~)f($\widetilde${V}), every edge is a geodesic segment between two vertices, and every face is isometric to a convex Euclidean polygon.
Denote the space of quasi-Fuchsian equivariant future-convex maps by $\widehat{\mathscr{E}}_{+}^{0}(V)$. By picking a lift in $\widetilde{V}$ for each $v\in V$, the space $\widehat{\mathscr{E}}_{+}(V)$ can be considered as a subset of $\widehat{\mathcal{R}}^{0}\times(\mathbb{R}^{2,1})^{V}$, from which it inherits a topology. This is an open subset, hence $\widehat{\mathscr{E}}_{+}^{0}(V)$ becomes a smooth manifold. Denote by $\mathscr{E}_{+}^{0}(V)$ its quotient by $G$. It is a smooth manifold of dimension 6k6+3n6k-6+3n, where $k$ is the genus of $S$ and $n:=|V|$. Similarly, we define quasi-Fuchsian equivariant past-convex maps and denote the respective spaces by $\widehat{\mathscr{E}}_{-}^{0}(V)$ and $\mathscr{E}_{-}^{0}(V)$. The notions of automorphic, trivial and isometric vector fields extend to this setting accordingly.
Let $\mathscr{S}^{0}(V)$ be the space of Euclidean cone-metrics with cone-angles $2\pi$ and vertices at $V$, considered up to isotopy. For $e\in\mathscr{E}_{+}^{0}(V)$ lifting to $(\rho,f)$, the induced metric on $\Sigma(f)$ can be associated with an element in $\mathscr{S}^{0}(V)$. This defines the induced metric map
@@
It was shown in [31, Lemma 2.13] that this map is $C^{1}$. Moreover, the authors proved
@@
This is a direct corollary of [31, Theorem I]. See also the discussion in [31, Section 4.1]. For $s\in\mathscr{S}^{0}(V)$ denote $(\mathscr{I}_{V,+}^{0})^{-1}(s)$ by $\mathscr{E}_{+}^{0}(s)$. Due to Theorem 5.6, this is a $C^{1}$-submanifold of $\mathscr{E}_{+}^{0}(V)$. We now define a map
@@
which sends a $\rho$-equivariant map to $\rho$ (recall that $\mathcal{R}^{0}\cong T\mathscr{T}$). It is also $C^{1}$, see [31, Section 4.1]. We define similarly the space $\mathscr{E}^{0}_{-}(s)$ and the map $\phi^{0}_{s,-}$. It is another direct corollary of [31, Theorem I] that
@@
For any Euclidean cone-metric $s$ with cone-angles $2\pi$, the compositions of $\phi_{s,\pm}^{0}$ with the projection to $\mathscr{T}$ are $C^{1}$-diffeomorphisms.
In particular, the maps $\phi^{0}_{s,\pm}$ are $C^{1}$-immersions. From [31, Theorem II], we have
@@
Pick two finite subsets $V_{+},V_{-}\subset S$ and metrics $s_{+}\in\mathscr{S}^{0}(V_{+})$, $s_{-}\in\mathscr{S}^{0}(V_{-})$. Then the intersection of $\phi^{0}_{s_{+},+}$ and $\phi^{0}_{s_{-},-}$ is unique and transverse.
@@
[…]
@@
We can now prove Proposition 5.4.
Proof of Proposition 5.4..
[…]
Recall the Pogorelov map $\Phi$ and the associated map $\Psi_{\rho}$. From Lemma 3.11, $\Phi(v)$ is a $\Psi_{\rho}(\dot{\rho})$-automorphic vector field. Also, if $v$ induces a zero variation on the distance between two points, then this means that its restriction to these two points coincides with the restriction of a Killing field. (Indeed, by subtracting a Killing field, one can assume that at one point $v$ is zero. Then at the second point $v$ is orthogonal to the segment between the points and any such $v$ is induced by an infinitesimal rotation around the first point.) Hence, Lemma 3.8 implies that $\Phi(v)$ is isometric. By Lemma 5.9, it is trivial, thereby so is $v$. It follows that $\dot{e}=0$. ∎